OpenStudy (thatonegirl_):

How do I figure this out? Suppose you had 2.0158grams of hydrogen (H2). A.How many moles of hydrogen do you have?(do i just divide it by the atomic weight?) B. How many moles of oxygen would react with this much hydrogen? C. What mass of oxygen would you need for this reaction? D. How many grams of water would you produce?

OpenStudy (aaronq):

A.How many moles of hydrogen do you have?(do i just divide it by the atomic weight?) yes, divide by the molar mass of \(H_2\) B to D: first write and balance the chemical equation for the process

OpenStudy (thatonegirl_):

when it asks how many moles would react with that much hydrogen, is hydrogen the product..? @aaronq

OpenStudy (aaronq):

nope, \(H_2 +O_2\rightarrow H_2O+ energy\) ill let you balance it

OpenStudy (thatonegirl_):

heres what i got a. 1 mole b.1 mole c. 32 amu d. 36.04 grams did i do that right? @aaronq

OpenStudy (aaronq):

you need to first balance the equation

OpenStudy (thatonegirl_):

yeah i did and i got\[2H _{2} O+O _{2} \rightarrow2H _{2} O\]

OpenStudy (thatonegirl_):

@aaronq

OpenStudy (aaronq):

you mean: \(2H_2+O_2→2H_2O\) so now, look at the coefficients, you used half as many oxygen molecules as hydrogen molecules in the reaction.

OpenStudy (thatonegirl_):

yea thats what i meant oops. @aaronq

OpenStudy (aaronq):

so now answer B-D

OpenStudy (thatonegirl_):

a) 2 moles..?? B) 1 mole c) 32 amu d) 36.04g @aaronq

OpenStudy (aaronq):

A was correct previously

OpenStudy (aaronq):

the rest are not

OpenStudy (thatonegirl_):

To figure out the moles, you divide the mass of hydrogen by 1.01 correct? @aaronq

OpenStudy (aaronq):

you divide by the molar mass of the molecule, which is dihydrogen \(H_2\), so the molar mass is 2*1.01=2.02

OpenStudy (thatonegirl_):

And am i dividing the 2.0158 grams or the 2H2O? @aaronq

OpenStudy (aaronq):

\(moles=\dfrac{mass}{molar~mass}\) \(moles_{H_2}=\dfrac{mass_{H_2}}{molar~mass_{H_2}}=\dfrac{2.0158 ~\cancel g}{2.0158 ~\cancel g/mol}=1~mole\)

OpenStudy (thatonegirl_):

oh okay. so how do i figure out moles of oxygen that react with that? @aaronq

OpenStudy (anonymous):

bro thats really simple you've got n=m/M you've got the mass and the molar which 1 you get n

OpenStudy (aaronq):

you use the coefficients from the balanced reaction \(\sf \dfrac{moles~of~H_2}{H_2's ~coefficient}=\dfrac{moles~of~O_2}{O_2's ~coefficient}\)

OpenStudy (aaronq):

set up that ratio with what you know (which is everything but moles of oxygen) and solve

OpenStudy (thatonegirl_):

1/2? @aaronq sorry my teacher doesn't teach -.-

OpenStudy (aaronq):

yep thats right

OpenStudy (aaronq):

now the mass of oxygen (\(O_2\)), comes from the same formula, rearranged: \(mass=moles*molar~mass\)

OpenStudy (thatonegirl_):

so is that 16?

OpenStudy (aaronq):

yes 16 grams

OpenStudy (aaronq):

to find the mass of water, you'd first find moles with a ratio (like the one earlier), then the same formula.

OpenStudy (thatonegirl_):

why wouldn't you just add the atomic weights of the molecule? @aaronq

OpenStudy (aaronq):

yeah, you can do that aswell (that's actually much much simpler)

OpenStudy (aaronq):

not atomic weights though, but masses

OpenStudy (thatonegirl_):

so 2x(2x1.01)+16?

OpenStudy (aaronq):

(2x1.01)+16 only

OpenStudy (thatonegirl_):

but in the balanced equations it's 2H2O?

OpenStudy (aaronq):

yeah but you only used 2 grams of \(H_2\) and 16 g of \(O_2\) the coefficients would only be used if you were finding moles first

OpenStudy (thatonegirl_):

ah okay. Thank you so much :D

OpenStudy (aaronq):

no problem !